Circular No. 6531

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996cd IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discovery

of an apparent supernova (V ∼ 18.5) on films taken on 1996 Dec. 16.06 and 16.12 UT. SN 1996cd is located at α = 7^{h}57^{m}20^{s}.73, δ = +11^{o}12′23′′.7 (equinox 2000.0), which is 15′′.5 east and 2′′.3 north of the center of a spiral galaxy (and located on the eastern arm of the galaxy). OCA films taken on 1997 Jan. 5.08 and 5.11 show the supernova about half a magnitude fainter than on the Dec. films. Nothing is visible at the position of SN 1996cd on 1995 OCA films or on the original Palomar Sky Survey prints. A nearby star (V ∼ 16) is at position end figures 21^{s}.07, 34′′.1.

SUPERNOVA 1997A IN ANONYMOUS GALAXY J. Maza, Department of Astronomy, University of Chile, reports the

discovery of a supernova on a CCD image taken on Jan. 3.32 UT with the Curtis-Schmidt telescope at Cerro Tololo Interamerican Observatory (CTIO), as part of a new Calan/Tololo supernova search involving P. F. Ortiz, L. Campusano, M. Wischnjewsky, and R. Antezana, University of Chile; M. M. Phillips, N. B. Suntzeff, R. Schommer, A. Clocchiatti, and R. Covarrubias, CTIO; and B. Schmidt, Mount Stromlo and Siding Spring Observatories. The supernova was confirmed via CCD images taken by M. T. Ruiz and M. Wischnjewsky at the 0.9-m CTIO telescope on Jan. 8.1, from which are estimated the following magnitudes: B = 18.52, V = 18.55. A position for SN 1997A measured from the Digital Sky Survey is α = 4^{h}31^{m}33^{s}.39, δ = −61^{o}07′10′′.4 (equinox 2000.0); the object is located 4′′.9 west and 0′′.1 north of the host galaxy’s center. SN 1997A is not present (to V = 20.5) on a CCD frame obtained by L. Campusano with the Curtis-Schmidt telescope on 1996 Dec. 14.2.

L. Wang, University of Texas; and D. Baade, European Southern Observatory (ESO), report their preliminary inspection of a spectrum (resolution ∼ 0.3 nm, range 385.5–801.7 nm) of SN 1997A obtained on Jan. 8.1 UT with the ESO 2.2-m telescope (+EFOSC2): “The spectrum clearly shows features of Si II, S II, and Fe III. The redshift of the host galaxy is ∼ 17 500 km/s, as measured from the Ca H and K lines. The expansion velocity, as measured from the minimum of the Si II feature, is ∼ 12 000 km/s after correcting for the redshift of the host galaxy. This is a type-Ia supernova prior to optical maximum.”

1997 January 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green